Show all work clearly. No credit for answer only. Provide the proper units. 1. A particle of mass 0.20 kg and charge 25.0 uC has an initial velocity of 20.0 m/s in the x-direction and 5.0 m/s in the positive y-direction. The particle enters a uniform electric field pointed downward in the negative y-direction. If the particles initial height above the ground is 15.0 meters, how far will the particle travel before it strikes the ground? E 2.0x10 NIC Note: Do not ignore gravity 6 1 First you need to find the acceleration and then the time Please explain how you get ( t ), which way you have used?. And why.


First you need to find the acceleration and then tFor this I only have one question. Do we need to divide the distance r 1 cm by 100 to convert it to m in order to get the final answer? Or just keep it as its.
Answer #1

2)

r has to be taken as 1/100 = 0.01 m. all Quantities must be in SI Units.

1)

F=q*E = m*a

Break this up along X and Y directions, along Y direction , Fy = qE=ma-mg

Calculate acceleration and use V= u+at to get time , t, then, s=ut+1/2 *at^2 to get distance